OpenStudy (anonymous):

Help Is really needed! I know I am reposting questions, but I'm really stuck and I need help.

OpenStudy (anonymous):

1. Which of the following equations could be the result of using the comparison method to solve the system shown? x + y = 5 2x + y = 7 2. For the following system, use the second equation to make a substitution for y in the first equation. 2x + y = 6 y = 3x + 4 What is the resulting equation? 3. Solve the following system of equations by the substitution method. 10x + 10y = 1 x = y - 3 What is the value of y? 4.Which of the following would not be the result of a substitution in the following system? 2x + y = 7 y - x = 1 5.Solve the following system of equations by the substitution method. x - y = 0 x - y - 2 = 0 What is the solution set? 6. Solve the following system of equations by the substitution method. 8x = 2y + 5 3x = y + 7 What is the solution set? 7.Based on the lesson, which of the following would be the best approach for solving this system by substitution? 5x = y + 6 2x - 3y = 4 8.Which of the following equations could be the result of using the comparison method to solve the system shown? x + 2y = 6 x - 4y = 8 9. Solve the following system of equations. 2x + y = 3 x = 2y - 1 10. For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 1 y + 4 = 5x What is the resulting equation? 11. For the following system, use the second equation to make a substitution for x in the first equation. x + 5y - 10 = 0 x = 2y - 8 What is the resulting equation in simplest form? 12. For the following system, use the second equation to make a substitution for x in the first equation. 3x + 2y = 7 x - y + 3 = 0 What is the resulting equation? 13. Solve the following system by the comparison method. 2p + q = 1 9p + 3q + 3 = 0 What is the solution set? 14.Which of the following equations could be the result of using the comparison method to solve the system shown? x - 4y - 1 = 0 x + 5y - 4 = 0 15.Solve the following system of equations. 3x + 2y - 5 = 0 x = y + 10